Please solve these factoring problems:
x2+6x+5
x2-8x+7
x2-4x+4
x2+x-42

Answer by stanbon(48574)

x2+6x+5 = (x+5)(x+1)
x2-8x+7 = (x-1)(x-7)
x2-4x+4 = (x-2)(x-2)
x2+x-42 = (x+7)(x-6)
-------------------------
Cheers,
Stan H.